OpenStudy (anonymous):

solve the following systems of equations by applying the method of Gauss: -2x-y+z=-4 3x+y-2z=6 2x+y+z=6

OpenStudy (aravindg):

@Muskan why do u find this difficult?

OpenStudy (anonymous):

i dont understand this equation the method of gauss

OpenStudy (aravindg):

solve 2 eqns to remove a variable ,solve the new eqn with the 3rd eqn

OpenStudy (anonymous):

I get no good answer

OpenStudy (helder_edwin):

u can also use a matrix and row-reduce it i think this is what u r supposed to do!

OpenStudy (aravindg):

ya u can indeed use matrices

OpenStudy (aravindg):

AX=B

OpenStudy (anonymous):

What I like to do when solving linear systems is to label my equations. For these lets label them from top to bottom, 1, 2, and 3. I see that if we add 1+2 or 1+3, we can eliminate a variable and create a new equation 4. Want to first give that a try?

OpenStudy (aravindg):

@agentc0re ME 2 :p

OpenStudy (anonymous):

i was tried..

OpenStudy (anonymous):

@AravindG It's the best way to keep track of things, don't ya think? :D great minds think alike.

OpenStudy (anonymous):

how can i solve ...

OpenStudy (anonymous):

Well what happens when you make two new equations by adding 1+2 and 1+3 like i suggested? call these new equations 4 and 5.

OpenStudy (aravindg):

u can solve when u make an honest attempt!!

OpenStudy (helder_edwin):

Come on @Muskan if we solve this for u it won't do u any good do what @agentc0re suggested and tell us what u got

OpenStudy (anonymous):

k..

OpenStudy (anonymous):

hey u have three equation and three unknown variable ! just find them a;;

OpenStudy (anonymous):

What did you get when you do what i suggested? Post your two new equations and we can go from there! :D

OpenStudy (aravindg):

@Muskan ?

OpenStudy (anonymous):

wait 2 min plz

OpenStudy (helder_edwin):

so?

OpenStudy (anonymous):

first equ: -2x-y+z=4 second equ: 3x+y-2z=6 ................................. -x -z=2 first equ: -2x-y+z=4 third equ: 2x+y+z=6 ............................. 2z=2 first equ: -2x-y+z=4 > -2*3-y+1=-4 > y=-4-1+6 > y=1 second equ: -x -z=2 > -x-1=2 > x=3 third equ: 2z=2 >z=2/2> z=1 y=1 x=3 z=1

OpenStudy (helder_edwin):

the sign of x is wrong

OpenStudy (anonymous):

-x=1+2> -x=3

OpenStudy (helder_edwin):

right -x=3 then x=-3

OpenStudy (anonymous):

okk

OpenStudy (helder_edwin):

u have to recompute the value of y

OpenStudy (anonymous):

but..still is bad..

OpenStudy (helder_edwin):

eres española? según tu perfil estudias en Linares

OpenStudy (anonymous):

si estoy en espana

OpenStudy (helder_edwin):

yo soy de Bolivia

OpenStudy (anonymous):

aa bien

OpenStudy (anonymous):

hablar ingles muy bien

OpenStudy (helder_edwin):

tu solución es x=-3, y=11 y z=1

OpenStudy (anonymous):

como y=11

OpenStudy (helder_edwin):

te dije que tenías que calcular de nuevo el valor de y

OpenStudy (anonymous):

first equ: -2x-y+z=4 > -2*3-y+1=-4 > y=-4-1+6 > y=1 yo hecho asi.. donde esta fallo

OpenStudy (helder_edwin):

tenías z=1 pero x=3 (que no era correcto) lo correcto ex z=1 y x=-3 entonces -2x-y+z=4 se convierte en -2(-3)-y+1=4 de donde y=4-1-6=-3

OpenStudy (anonymous):

pero no es y=11 es y=-3

OpenStudy (helder_edwin):

lo siento cometí un error z=1 x=3 y=-1

OpenStudy (anonymous):

ahora z=1 y=-3 x=-3 es correcto

OpenStudy (helder_edwin):

en lo que escribiste arriba -x-z=2 el signo de x es incorrecto debería ser x-z=2 de donde x=3 (pues z=1)

OpenStudy (anonymous):

aah vale

OpenStudy (helder_edwin):

entonces -2x-y+z=-4 se convierte en -2(3)-y+1=-4 de donde -y=-4-1+6=1 de donde y=-1

OpenStudy (anonymous):

man.wat language is this!!!!!! strange!

OpenStudy (anonymous):

yuhuuuuuuuuuuuuu I gottttt it....

OpenStudy (anonymous):

@Yahoo! is spanish

OpenStudy (anonymous):

u r frm Spain

OpenStudy (anonymous):

i thought u are frm PAK

OpenStudy (anonymous):

m 4rom pak but i live in spain

OpenStudy (anonymous):

oh.......thats Fine!

OpenStudy (anonymous):

m happy i got my ans...

OpenStudy (anonymous):

i think u have done well...))

OpenStudy (helder_edwin):

buen trabajo @Muskan !

OpenStudy (anonymous):

gracias :))

OpenStudy (helder_edwin):

Adios, y cuídate

OpenStudy (anonymous):

hasta luego..

OpenStudy (anonymous):

igualmente